OpenStudy (anonymous):

Simplify 6 4*√4 - 7 *3√4 + 2 *3√4 - 3 *3√4.

OpenStudy (anonymous):

64*√4 - 7 *3√4 + 2 *3√4 - 3 *3√4. or 6*4*√4 - 7 *3√4 + 2 *3√4 - 3 *3√4 ?

OpenStudy (anonymous):

lol yes i find this confusin also lol

OpenStudy (anonymous):

the 6 and the 4 are separate. it's 6 4*√4 - 7 *3√4 + 2 *3√4 - 3 *3√4. or 6*4*√4 - 7 *3√4 + 2 *3√4 - 3 *3√4 by the asteric i mean that the number before it is a small raised number.

OpenStudy (anonymous):

so the numers asterick are cube/4th roots yada yada?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

so in other words it is this \[6\sqrt[4]{4}\]

OpenStudy (anonymous):

as the first part?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

alright i'll equate it out for nick so he can make sure i'm sure he's got hte right answer \[6\sqrt[4]{4}-8\sqrt[3]{4}+2\sqrt[3]{4}-3\sqrt[3]{4}\]

OpenStudy (anonymous):

6*4√4 - 7 *3√4 + 2 *3√4 - 3 *3√4 3√4(6*√4-7+2-3) 3√4(6*√4-8)

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

yes. (:

OpenStudy (anonymous):

nick's answer is right or the way i showed it/

OpenStudy (anonymous):

out u wrote 8 instead of 7

OpenStudy (anonymous):

and 95 girl i am right i just simplified it

OpenStudy (anonymous):

i took 3√4 common

OpenStudy (anonymous):

out u do it too y will get the same one wht i got

OpenStudy (anonymous):

it's correct it just looks so weird when you use a check to show sqroot i see it as all under the root

OpenStudy (anonymous):

its correct so i deserve a medal too bro

OpenStudy (anonymous):

lol